# Latent Diffusion Models for Controllable RNA Sequence Generation

### v1

@@ Abstract
This paper presents RNAdiffusion , a latent diffusion model for generating and optimizing discrete RNA sequences. RNA is a particularly dynamic and versatile molecule in biological processes. RNA sequences exhibit high variability and diversity, characterized by their variable lengths, flexible three-dimensional structures, and diverse functions. We utilize pretrained BERT-type models to encode raw RNAs into token-level biologically meaningful representations. A Q-Former is employed to compress these representations into a fixed-length set of latent vectors, with an autoregressive decoder trained to reconstruct RNA sequences from these latent variables. We then develop a continuous diffusion model within this latent space. To enable optimization, we train reward networks to estimate functional properties of RNA from the latent variables. We employ gradient-based guidance during the backward diffusion process, aiming to generate RNA sequences that are optimized for higher rewards. Empirical experiments confirm that RNAdiffusion generates non-coding RNAs that align with natural distributions across various biological indicators. We fine-tuned the diffusion model on untranslated regions (UTRs) of mRNA and optimize sample sequences for protein translation efficiencies. Our guided diffusion model effectively generates diverse UTR sequences with high Mean Ribosome Loading (MRL) and Translation Efficiency (TE), surpassing baselines. These results hold promise for studies on RNA sequence-function relationships, protein synthesis, and enhancing therapeutic RNA design.
@@ 1 Introduction
Beyond continuous domains, people have explored using diffusion models to generate discrete sequences to inherit the benefits of controllability through guidance, both in the textual domain [14, 77] and in the biology domain [37, 126, 88, 45]. We note that in the domain of text generation, state-of-the-art performances are achieved by scaling up auto-regressive models [2, 10, 116, 118, 119]. However, for generating biological sequences such as proteins, DNAs, and RNAs, language models have to be adapted to handle their specific sequence characteristics [101, 65, 94, 19, 63, 22].
In this paper, we explore generative models for RNAs, which introduce distinct challenges in machine learning due to their inherent complexity. RNA sequences demonstrate high variability and diversity. Several portions of the RNA sequence may be missing and vary dramatically in length [101] due to processes like alternative splicing [89], which allows splicing a gene sequence in multiple ways to produce different isoforms. Additionally, unlike proteins, the three-dimensional structures of many RNAs remain unresolved [120], further complicating the modeling of their functions. Generative models for RNA are particularly useful for designing untranslated regions (UTR) of mRNA vaccines, which plays a central role in regulating gene expression [92, 25, 13] and mRNA stability [12]. The ability to generate optimized UTR sequences offers significant potential for controlling gene expression and advancing targeted therapies [103, 24, 18]. Thus, we aim to develop a robust, powerful language model for generating non-coding RNAs and optimizing UTRs.
@@
Figure 1: RNAdiffusion: Latent diffusion model for RNA sequences. It has three parts: (1) RNA sequence auto-encoder, consisting of a pretrained RNA-FM model and a Querying Transformer, for translating between the sequence space and the latent space; (2) Guided diffusion model with a pre-trained score network, for generating latent RNA embeddings under external guidance; (3) Latent reward model, trained on the latent space to predict functional properties of RNA, for computing guidance of diffusion.
@@
The functional properties of RNAs, especially non-coding RNAs, involve complex, nonlinear interactions within the sequence [95], which do not conform to a simple left-to-right progression. Thus, autoregressive models are inadequate for capturing their patterns and dynamics. In this paper, we turn to diffusion models, which allow for the holistic modeling and generation of RNA sequences. There are several technical challenges. Individual nucleotides (A, U, G, C) do not exhibit meaningful biological functions independently. Instead, these nucleotides form higher-order structures, such as motifs, which contribute to the functional properties of ncRNAs. Thus, the denoising network has to account for both the high-level picture of the entire sequence and the low-level syntax. Additionally, RNAs have dramatically different sequence lengths, adding difficulty to efficient training and generalization. Although Transformer denoisers can handle variable-length inputs in theory, people still either pad/truncate input to a fixed length or build an additional length prediction module to ensure robust practical performances [105, 64, 130].
@@
To overcome these challenges, we develop diffusion models that operate on a latent space for RNA sequences. Inspired by recent advances [102, 87], our approach utilizes a sequence auto-encoder to map raw input sequences into a latent embedding space. We then train a score-based diffusion model to learn this latent distribution. Refer to Figure 1 for a schematic of our pipeline. This latent-space approach is particularly effective for RNA sequences for several reasons. First, it delegates the generation of the intricate details of raw RNA sequences to the decoder, allowing the diffusion model to concentrate on capturing the coarse-grained distribution in the meaningful latent space. Second, the latent space is designed to be of fixed size, simplifying both the design and training of the diffusion model by eliminating the need for additional modules or adaptations to accommodate varying sequence lengths.
@@
For the sequence autoencoder, we adopt a pretrained RNA language model RNA-FM [26] to map raw sequences to biologically-meaningful token-level representations. Then we use a trainable Q-Former [76] to summarize the token representations into a fixed-length list of latent vectors. Next, we train a standard score-based denoising diffusion model to model the distribution of RNAs in the latent space. For validation, we analyzed the set of generated sequences and examined a range of biological indicators, such as minimal Levenshtein distance, minimal 4-mer distance, G/C content, minimum free energy, secondary structure, etc. Empirical experiments confirm that the generated sequences align with natural distributions of non-coding RNAs across these biological indicators. In other words, our RNA latent diffusion models can generate novel RNA sequences that are highly similar to natural sequences.
@@
As latent variables from Q-Formers provide biologically-meaningful summarization, we opt to train reward networks directly in the latent space to predict RNA’s functional properties. Our rewards of interests are the Mean Ribosome Loading (MRL) [103] and the Translation Efficiency (TE) [24], which measure the mRNA-to-protein production levels and efficiencies. These functional properties of mRNA are largely determined by UTRs. By using trained reward models, we compute their gradients and construct guidance for the diffusion model. Thus, we can generate optimized UTR sequences toward higher values of MRL and TE by adding this guidance to the backward generation process of our diffusion model, which is fine-tuned upon a ncRNA-pretrained model. Adding guidance is done in a plug-and-play manner, without additional training. Empirically, the generated UTRs from guided diffusion achieve a 166.7% improvement of TE and an 52.6% improvement of MRL compared to mRNA baselines.
We summarize our contributions below:
  • We develop the first latent diffusion model for RNAs. Our Sequence AutoEncoder involves a novel adaptation of Querying-Transformers that summarizes the varying-length output from pretrained encoders into fixed-length embeddings. Empirical study confirmed that our model can generate non-coding RNAs that align with natural distributions across various biological indicators.
  • The latent diffusion model allows controlled generation via guidance. We demonstrated optimization of 5’-UTR sequences towards high protein production levels, where our result improves the TE by up to $166.7\%$ and MRL by up to $52.6\%$.
@@ 2.1 Sequence Auto-Encoder
Querying-Transformer. Next we use a Q-Former, denoted by $\mathrm{QFormer}_{\phi}(\cdot)$, to summarize the embeddings of the pre-trained encoder into $K$ latent vectors $\bm{z}=(\bm{z}^{1},\bm{z}^{2},\dots,\bm{z}^{K})$, where $K$ is a fixed number independent of $L$. We aim to achieve three goals: (1) To retains enough information in latent vectors, making it easy to train a decoder for raw sequence reconstruction; (2) To use latent vectors later for predicting RNA functional properties, via training additional prediction models on top of these embedding; (3) To reduce the dimensionality of raw sequences, so that diffusion models can focus on learning intrinsic structure of latent RNA vectors.
The Querying Transformer (Q-Former) [76] takes $K$ trainable query token embeddings $(\bm{q}^{1},\dots,\bm{q}^{K})$ as input. The query tokens go through the transformer with cross-attentions to the embedding sequence given by the encoder and progressively summarize the original token embeddings of varying lengths into a fixed-size sequence of embedding vectors. This is a flexible approach compared to using only the <cls> token embedding  [75]. As when we increase $K$, we retain more information in the latent vectors and allow accurate reconstruction, which helps to mitigate the KL-vanishing problem [21].
@@
$\mathcal{L}_{\text{reconstruct}}(\psi,\phi;\bm{x})=-\log p_{{\mathrm{Dec}}_{\psi}}(\bm{x}^{1},\bm{x}^{2},\dots,\bm{x}^{L}\mid\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))),$
@@ 2.2 Latent Diffusion Model
$\mathcal{L}(\theta)=\mathbb{E}_{\bm{z}_{0}=\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))}\mathbb{E}_{t,\bm{\epsilon}}\|\bm{\epsilon}-\mathrm{Denoiser}_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\|^{2},$
In our framework, the latent embedding have fixed sizes, making it easy to train diffusion models and treating data as if they are continuous like images. Unlike other discrete diffusion models, our framework does not require truncating/padding sequences to a fixed length compared with [77, 14], and it does not require training an additional length predictor as done by  [105, 64, 130]. Our auto-regressive decoder has been trained to generate sequences with variable lengths.
@@ 2.3 Guided Generation in the Latent Space
$\mathcal{L}(\xi)=\operatorname{\mathbb{E}}_{(\bm{x},r)}\ell(R_{\xi}(\bm{z}),r),$
where $\ell$ is a loss function, $\bm{z}=\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))$ is the embedding obtained from Q-Transformer, and $\xi$ are weights of the reward model.
@@
$\widehat{\bm{z}}_{0}=\widehat{\bm{z}}_{0}(\bm{z}_{t})=\frac{\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathrm{Denoiser}_{\theta}(\bm{z}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$
@@
predicted $\text{predicted noise}=\mathrm{Denoiser}_{\theta}(\bm{z}_{t},t)+\sqrt{1-\bar{\alpha}_{t}}\cdot\lambda\cdot\nabla_{\bm{z}_{t}}\ell(r^{*},R_{\xi}(\widehat{\bm{z}}_{0})).$
@@ 3.1 Experiment Setup
Models. We adopt the pretrained RNA-FM [26] as the encoder. RNA-FM is an encoder-only model trained with masked language modelling on non-coding RNAs (ncRNAs), which can extract biologically meaningful embeddings from ncRNAs and can be adapted for various downstream tasks, such as UTR function prediction, RNA-protein interaction prediction, secondary structure prediction, and etc. For the Q-Former, we modify the ESM-2 model architecture [82] by adding $K$ trainable query token embeddings and adding cross-attention layers in each transformer block that attends to the embeddings from the RNA-FM. For the decoder model, we adopt the same model architecture of ProGen2-small [94] with our customized tokenizers for RNAs.
Dataset. We pretrain our RNA sequence Auto-encoder with the ncRNA subset of Ensembl database [90] downloaded from RNAcentral [1], which contains collected ncRNA sequences from a variety of vertebrate genomes and model organisms. We select the sequences shorter than 768 bp to avoid exceeding RNA-FM’s maximum length limit and exclude those with rare tokens more than {A, C, G, U}, curating a dataset of 1.1M samples for training and testing. Due to the space limit, additional experimental setup can be found in Appendix D.1.
@@ 3.2 Results
We perform a sweep over hyperparameters and obtain a set of RNA Sequence AutoEncoders with high reconstruction capabilities. Specifically, there are two hyperparameters that influence the dimension of the latent space: the number of query tokens $K$ and the hidden dimension $D$ of each latent vector. We tune $K$ among $\{16,32\}$ and $D$ among $\{40,80,160,320\}$. For each input sequence $\bm{x}$ of length $L$, we calculate the length-normalized reconstruction loss (NLL), i.e., $-\frac{1}{L}\log p_{\mathrm{Dec}_{\psi}}(\bm{x}\mid\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))).$ We also calculate the length-normalized edit distance (NED) between the reconstructed sequence $\widehat{\bm{x}}\sim p_{\mathrm{Dec}_{\psi}}(\cdot\mid\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x})))$ and the original input sequence $\bm{x}$, i.e., $\frac{1}{L}\mathrm{EditDistance}(\widehat{\bm{x}},\bm{x}).$ For each of the two metrics, we report the empirical average on a held-out test set in Table 1. We see that increasing either the number of query tokens $K$ or the hidden dimension $D$ reduces the reconstruction errors and the error rates. Increasing the dimension of the latent space retains more information in the latent embeddings, and hence it is easier to reconstruct the input sequence.
Table 1: The performance of the RNA sequence auto-encoders and the latent diffusion models. The ELBO is calculated using Equation (12) of [57].
Query Token Length $K$ Embedding Dim $D$ Sequence AutoEncoder Latent Diffusion
Reconstruction Error (NLL)$\downarrow$ Normalized Edit Distance $\downarrow$ NLL (bits/dim)$\downarrow$
@@
4 Latent Diffusion Model for RNAs

### v2

@@ Abstract
This work presents RNAdiffusion , a latent diffusion model for generating and optimizing discrete RNA sequences of variable lengths. RNA is a key intermediary between DNA and protein, exhibiting high sequence diversity and complex three-dimensional structures to support a wide range of functions. We utilize pretrained BERT-type models to encode raw RNA sequences into token-level, biologically meaningful representations. A Query Transformer is employed to compress such representations into a set of fixed-length latent vectors, with an autoregressive decoder trained to reconstruct RNA sequences from these latent variables. We then develop a continuous diffusion model within this latent space. To enable optimization, we integrate the gradients of reward models—surrogates for RNA functional properties—into the backward diffusion process, thereby generating RNAs with high reward scores. Empirical results confirm that RNAdiffusion generates non-coding RNAs that align with natural distributions across various biological metrics. Further, we fine-tune the diffusion model on mRNA 5’ untranslated regions (5’-UTRs) and optimize sequences for high translation efficiencies. Our guided diffusion model effectively generates diverse 5’-UTRs with high Mean Ribosome Loading (MRL) and Translation Efficiency (TE), outperforming baselines in balancing rewards and structural stability trade-off. Our findings hold potential for advancing RNA sequence-function research and therapeutic RNA design.
@@ 1 Introduction
While diffusion models have been widely applied in continuous domains, they have also proven useful in generating discrete sequences, inheriting the benefits of controllability through guidance. This has been shown both in the textual domain [14, 77] and in the biology domain [37, 122, 88, 45]. In the domain of text generation, state-of-the-art performances are achieved by scaling up auto-regressive models [2, 10, 113, 115, 116]. However, for generating biological sequences such as proteins, DNA, and RNA, language models have to be adapted to handle their specific sequence characteristics [99, 65, 93, 19, 63, 22].
In this work, we focus on developing generative models for RNA, with applications to both non-coding RNAs (ncRNAs) and mRNA untranslated regions (UTRs).
For ncRNAs, they are crucial regulators of gene expression and cellular processes, influencing pathways involved in development, disease, and immune responses. In parallel, the UTRs of mRNAs play a central role in regulating gene expression [91, 25, 13] and mRNA stability [12]. The ability to generate optimized UTR sequences offers significant potential for controlling gene expression, advancing mRNA vaccine designs, and enabling targeted therapies [101, 24, 18]. Therefore, our goal is to develop a robust and versatile language model capable of generating ncRNAs, and optimizing UTRs for enhanced functionality.
@@
Figure 1: RNAdiffusion : Latent diffusion model for RNA sequences. Three parts of RNAdiffusion : (1) RNA sequence auto-encoder, consisting of a pretrained RNA-FM model, a Querying Transformer, and a decoder, for translating between the sequence space and the latent space; (2) Guided diffusion model with a pre-trained score network, for generating latent RNA embeddings under external guidance; (3) Latent reward model, trained on the latent space to predict functional properties of RNA, for computing guidance of diffusion.
@@
For ncRNAs and UTRs, individual nucleotides (A, U, G, C) often do not independently exhibit functions. Instead, the nucleotides may form clusters of higher-order structures, such as motifs, contributing to the functional properties of the RNA. This introduces distinct challenges in sequence modeling. Specifically, the denoising network has to account for the high-level picture of the entire sequence and the low-level syntax within a local cluster. Besides, these RNAs have dramatically different lengths, adding difficulty to efficient training and generalization. Although Transformer denoisers can handle variable-length inputs in theory, existing studies mostly pad or truncate input to a fixed length, or build an additional length prediction module to ensure robust practical performances [103, 64, 126].
@@
To overcome these challenges, we develop diffusion models on a latent space for RNA sequences. Inspired by recent advances [100, 87], our approach utilizes a sequence auto-encoder to map raw input sequences into a latent embedding space. We then train a score-based continuous diffusion model to learn this latent distribution. Please see Figure 1 for a detailed schematic. This latent-space approach is effective for RNA sequences: First, it delegates the generation of the intricate raw RNA sequences to the sequence decoder, allowing the diffusion model to concentrate on capturing the coarse-grained distribution in the meaningful latent space. Second, the latent space is of fixed size, simplifying the design and training of the diffusion model by eliminating the need for additional modules or special adaptations for varying sequence lengths.
@@
For the sequence autoencoder, we use a pretrained language model RNA-FM [26] to map raw sequences to meaningful, token-level representations. A trainable Query Transformer (Q-Former) [76] summarizes these representations into fixed-length latent vectors. Next, we train a score-based denoising diffusion model to capture the distribution of RNAs in latent space. We then assess the generated ncRNAs using biological metrics such as Levenshtein distance, 4-mer distance, G/C content, minimum free energy, and secondary structure. Results confirm that the generated ncRNAs align with natural ncRNA distributions across these indicators. Thus, our latent diffusion model can produce novel RNA sequences with high similarity to native ncRNAs.
@@
Since latent variables from Q-Formers provide biologically meaningful summaries, we train reward networks directly in the latent space to predict functional properties of UTRs, key elements in mRNAs. The rewards, Mean Ribosome Loading (MRL) [101] and Translation Efficiency (TE) [24], measure mRNA-to-protein production levels and efficiencies, respectively. Using the trained reward models, we compute gradients and guide the diffusion model. Results show that RNAdiffusion generates novel UTRs with substantially higher MRL and TE values by incorporating guidance into the backward generation process. This guidance is a plug-and-play method that requires no extra training. Overall, the generated UTRs from guided diffusion achieve a 166.7% improvement of TE and a 52.6% improvement of MRL compared to unguided generation.
We summarize the contributions of RNAdiffusion :
  • Our sequence autoencoder, a novel adaptation of Q-Former, summarizes varying-length outputs from pretrained encoders into fixed-length embeddings, accurately reconstructing input sequences.
  • Our latent diffusion model generates novel RNA sequences, and empirical studies confirm that these sequences align with natural distributions across various biological metrics.
  • RNAdiffusion enables controlled generation via guidance, optimizing 5’-UTRs for higher protein production ($166.7\%$ improved TE and $52.6\%$ improved MRL compared to unguided generation), while achieving a better reward–structural stability trade-off than baselines.
@@ 2.1 Sequence Auto-Encoder
Querying-Transformer. Next we use a Q-Former, denoted by $\mathrm{QFormer}_{\phi}(\cdot)$, to summarize the embeddings of the pre-trained encoder $\mathrm{Enc}(x)$ into $K$ latent vectors $\bm{z}=(\bm{z}^{1},\bm{z}^{2},\dots,\bm{z}^{K})$, where $K$ is a fixed number independent of $L$. We aim to achieve three goals: (1) To retains enough information in latent vectors, making it easy to train a decoder for raw sequence reconstruction; (2) To use latent vectors later for predicting RNA functional properties, via training additional prediction models on top of these embedding; (3) To reduce the dimensionality of raw sequences, so that diffusion models can focus on learning intrinsic structure of latent RNA vectors.
The Querying Transformer (Q-Former) [76] takes $K$ trainable query token embeddings $(\bm{q}^{1},\dots,\bm{q}^{K})$ as input. The query tokens go through the transformer with cross-attentions to the embedding sequence given by the encoder and progressively summarize the original token embeddings of varying lengths into a fixed-size sequence of embedding vectors. This is a flexible approach compared to using only the <cls> token embedding  [75]. As when we increase $K$, we retain more information in the latent vectors and allow accurate reconstruction, which helps to mitigate the KL-vanishing problem [21]. Our ablation study in Section 3.3 validates the necessity of the Q-former.
@@
$\mathcal{L}_{\text{reconstruct}}(\psi,\phi;\bm{x})=-\log p_{{\mathrm{Dec}}_{\psi}}(\bm{x}^{1},\bm{x}^{2},\dots,\bm{x}^{L}\mid\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))),$ (1)
@@ 2.2 Latent Diffusion Model
$\mathcal{L}(\theta)=\mathbb{E}_{\bm{z}_{0}=\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))}\mathbb{E}_{t,\bm{\epsilon}}\|\bm{\epsilon}-\mathrm{Denoiser}_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},t)\|^{2},$ (2)
In our framework, the latent embedding have fixed sizes, making it easy to train diffusion models and treating data as if they are continuous like images. Unlike other discrete diffusion models, our framework does not require truncating/padding sequences to a fixed length compared with [77, 14], and it does not require training an additional length predictor as done by  [103, 64, 126]. Our auto-regressive decoder has been trained to generate sequences with variable lengths; see Figure 2.
@@ 2.3 Guided Generation in the Latent Space
$\mathcal{L}(\xi)=\operatorname{\mathbb{E}}_{(\bm{x},r)}\ell(R_{\xi}(\bm{z}),r),$ (3)
where $\ell$ is a loss function, $\bm{z}=\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x}))$ is the embedding obtained from Q-Former, and $\xi$ are weights of the reward model.
@@
$\widehat{\bm{z}}_{0}=\widehat{\bm{z}}_{0}(\bm{z}_{t})=\frac{\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\mathrm{Denoiser}_{\theta}(\bm{z}_{t},t)}{\sqrt{\bar{\alpha}_{t}}}.$ (4)
@@
predicted $\text{predicted noise}=\mathrm{Denoiser}_{\theta}(\bm{z}_{t},t)+\sqrt{1-\bar{\alpha}_{t}}\cdot\lambda\cdot\nabla_{\bm{z}_{t}}\ell(r^{*},R_{\xi}(\widehat{\bm{z}}_{0})).$ (5)
@@ 3.1 Experiment Setup
Models. We adopt the pretrained RNA-FM [26] as the encoder. RNA-FM is an encoder-only model trained with masked language modelling on non-coding RNAs (ncRNAs). The architecture of our Q-Former is modified from the ESM-2 model structure [82]. For the decoder model, we adopt the same model architecture of ProGen2-small [93] with our customized tokenizers for RNAs.
Dataset. We pretrain our RNA sequence autoencoder with the ncRNA subset of Ensembl database [89] downloaded from RNAcentral [1], which contains collected ncRNA sequences from a variety of vertebrate genomes and model organisms. We select the sequences shorter than 768 bp to avoid exceeding RNA-FM’s maximum length limit and exclude those with rare tokens more than {A, C, G, U}, curating a dataset of 1.1M samples for training and testing.
Hyperparameter Search. There are two hyperparameters that influence the dimension of the latent space: the number of query tokens $K$ and the hidden dimension $D$ of each latent vector. We tune $K$ among $\{16,32\}$ and $D$ among $\{40,80,160,320\}$.
Evaluation Metrics. For each input sequence $\bm{x}$ of length $L$, we calculate two metrics: (1)length-normalized reconstruction loss (i.e., negative log-likelihood, NLL), i.e., $-\frac{1}{L}\log p_{\mathrm{Dec}_{\psi}}(\bm{x}\mid\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x})));$ (2)length-normalized edit distance (NED) between the reconstructed sequence $\widehat{\bm{x}}\sim p_{\mathrm{Dec}_{\psi}}(\cdot\mid\mathrm{QFormer}_{\phi}(\mathrm{Enc}(\bm{x})))$ and the original input sequence $\bm{x}$, i.e., $\frac{1}{L}\mathrm{EditDistance}(\widehat{\bm{x}},\bm{x}).$ Due to the space limit, additional experimental setup can be found in Appendix D.1.
@@ 3.2 Results
We perform a sweep over hyperparameters and obtain a set of RNA sequence autoencoders with high reconstruction capabilities. For each of the two metrics, we report the empirical average on a held-out test set in Table 1. We see that increasing either the number of query tokens $K$ or the hidden dimension $D$ reduces the reconstruction errors and the error rates. Increasing the dimension of the latent space retains more information in the latent embeddings, and hence it is easier to reconstruct the input sequence.
Table 1: The performance of the RNA sequence autoencoders and the latent diffusion models. For diffusion models, the upper bound of NLL, i.e., ELBO, is calculated using Equation (12) of [57].
$K$ $D$ Sequence AutoEncoder Latent Diffusion
[…]
@@
3.3 Ablation Studies.
We perform ablation studies to validate the necessities of the components of our sequence auto-encoder. When we replace the Q-former with a naive average pooling of RNA-FM output token embeddings $\mathrm{Enc}(x)$, the reconstruction performance of the Sequence AutoEncoder is significantly dampened, resulting in a 34.8% normalized edit distance (see Table 2).
We also experiment with replacing the RNA-FM with naive one-hot token embeddings and the results indicate that both the reconstruction capability of the sequence autoencoder and the reward modeling ability of the latent vectors are negatively affected, which validate the necessity of using the pre-trained RNA-FM as the encoder. Full results are deferred to Appendix D.4.
Settings ($K\times{D}$) Sequence AutoEncoder
NLL $\downarrow$ NED $\downarrow$
w/o Q-former ($1\times 640$) 1.0061 34.78 % $\pm$ 17.81 %
w/ Q-former ($16\times 160$) 0.0156 2.19% $\pm$ 5.50 %
4 Latent Diffusion Model for ncRNAs